ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

20. Special Activities of the ICM 1998

1991 MS Classification: 11, 68 Conrad, Marc, Universität des Saarlandes, Germany SIMATH SIMATH is a computer algebra system, focusing on computational number theory and especially on the arithmetic of elliptic curves. SIMATH contains a comprehensive library of C-functions and the calculator simcalc which is a user interface providing easy access to the library functions. Because of simcalc SIMATH is well suited for two different user groups. On the one hand, the calculator simcalc can be used for quick interactive calculations, where the user needs not to care about type declarations or other technical details with respect to programming languages. On the other hand, the experienced C or C++ programmer may use the SIMATH C-libraries to develop his own programs in the context of advanced calculations in number theory. See also http://emmy.math.uni-sb.de/∼simath for further details.